1) The magnitude of the Earth’s magnetic field is about 0.5 Gauss near Earth's surface. What’s the maximum possible force on an electron near the Earth’s surface with kinetic energy of 1 keV? How does this compare to the gravitational force on the electron near the Earth’s surface?


1
Expert's answer
2021-03-03T11:24:29-0500

The maximum modulus of the Lorentz force is "F_L = q V B". We may obtain velocity from the kinetic energy.

"E = \dfrac{m_eV^2}{2}, \;\; V = \sqrt{\dfrac{2E}{m_e}} = \sqrt{\dfrac{2\cdot 10^3\cdot1.6\cdot10^{-19}\,\mathrm{J}}{9.1\cdot10^{-31}\,\mathrm{kg}}} = 1.9\cdot10^7\,\mathrm{m/s}"

The modulus of the Lorentz force is "F_L = 1.6\cdot10^{-19}\,\mathrm{C}\cdot 1.9\cdot10^{7}\,\mathrm{m/s}\cdot 0.5\cdot10^{-4}\,\mathrm{T}= 1.52\cdot 10^{-16}\,\mathrm{N}."

The gravitational force is "F_G = \dfrac{GM_{\oplus}m_e}{R_{\oplus}^2} = m_e\cdot g = 9.1\cdot10^{-31}\,\mathrm{kg}\cdot 9.81\,\mathrm{N/kg} = 8.9\cdot10^{-30}\,\mathrm{N}."

We can see that the Lorentz force is significantly greater than the gravitational force.